# LaB-RAG: Label Boosted Retrieval Augmented Generation for Radiology Report Generation

### v1

@@ Abstract
In the current paradigm of image captioning, deep learning models are trained to generate text from image embeddings of latent features. We challenge the assumption that these latent features ought to be high-dimensional vectors which require model fine tuning to handle. Here we propose Label Boosted Retrieval Augmented Generation (LaB-RAG), a text-based approach to image captioning that leverages image descriptors in the form of categorical labels to boost standard retrieval augmented generation (RAG) with pretrained large language models (LLMs). We study our method in the context of radiology report generation (RRG), where the task is to generate a clinician’s report detailing their observations from a set of radiological images, such as X-rays. We argue that simple linear classifiers over extracted image embeddings can effectively transform X-rays into text-space as radiology-specific labels. In combination with standard RAG, we show that these derived text labels can be used with general-domain LLMs to generate radiology reports. Without ever training our generative language model or image feature encoder models, and without ever directly “showing” the LLM an X-ray, we demonstrate that LaB-RAG achieves better results across natural language and radiology language metrics compared with other retrieval-based RRG methods, while attaining competitive results compared to other fine-tuned vision-language RRG models. We further present results of our experiments with various components of LaB-RAG to better understand our method. Finally, we critique the use of a popular RRG metric, arguing it is possible to artificially inflate its results without true data-leakage. Our code is available at: https://github.com/uc-cdis/label-boosted-RAG-for-RRG.
11footnotetext: These authors contributed equally22footnotetext: Corresponding author: rgrossman1@uchicago.edu11footnotetext: Department of Computer Science, University of Chicago22footnotetext: Medical Scientist Training Program, Pritzker School of Medicine33footnotetext: Center for Translational Data Science, University of Chicago44footnotetext: Department of Medicine, University of Chicago00footnotetext: Preprint. Under Review.
@@ 1 Introduction
Figure 1: Overview of LaB-RAG for RRG.
Radiology reports are free-text natural language notes describing the observations seen in radiological images, such as X-rays, CT scans, or MRI scans. These reports are written by board-certified radiologists, doctors who have had years of medical training to read these imaging modalities. This is a critically important task as radiological imaging provides a noninvasive view of internal organs and tissues, offering clinicians insights into how best to intervene and treat disease. However, the skills required to read these images can be highly specialized [67]; there is also an increasing demand for radiologists to read the growing number of images collected [39, 13, 66]. While remote radiology increased with the COVID-19 pandemic, these services still face challenges affecting the limited radiologist workforce [26]. Motivated by these considerations and coupled with the popularization of large language models (LLMs), there has been an increasing interest in developing AI tools to help bridge the gap for the radiologist shortage [24, 52, 38].
@@
Radiology report generation (RRG) is the task of automatically generating these reports given the image [69, 47]. While RRG can be applied in any radiological imaging modality, the most frequent modality studied in the literature is X-rays of the chest. This is likely due to the frequent and inexpensive collection of chest X-rays (CXRs) in the clinical setting and the public availability of paired chest X-ray and report datasets [69], such as MIMIC-CXR [35, 36], OpenI [14], PadChest [7], BIMCV-COVID19 [77], VinDr-CXR [55], and others. For CXRs, the task is typically formulated as generating the “Findings”, “Impression”, or both sections of the corresponding report [47]. Conceptually, the findings section describes all positive or negative observations seen in the X-ray, while the impression section summarizes and interprets those findings with recommendations for clinical diagnosis [59].
At its core, RRG is a form of image captioning, specialized for medical imaging [71]. In recent years, LLMs have demonstrated impressive performance across the medical domain, passing and excelling at US medical board exams and outperforming clinicians in certain tasks [58, 11, 68]. While these results do not translate directly to clinical viability, medical image captioning is a natural area of research application for such LLMs. Efforts typically involve fine-tuning vision-language models for the medical image modality of interest [3, 1]. Even when starting with a pre-trained language, vision, or multimodal foundation models (FMs), domain adaptation is typically expected when the task is specific and potentially out-of-distribution [6].
@@
Model adaptation classically involves model training via supervised fine tuning (SFT). SFT of FMs is becoming increasingly difficult as models become larger, requiring compute resources greater than consumer-grade workstations can provide [75]. While parameter-efficient fine tuning (PEFT) methods have demonstrated competitive results [16], a form of LLM adaptation that does not require model training is in-context learning (ICL) [17]. The goal of ICL is to have an LLM infer the target output using examples of input-output pairs given jointly at inference time with the target input [49]. Related to ICL is retrieval augmented generation (RAG), a framework for providing additional context to an LLM prompt by retrieving documents related to the input query [42]. ICL and RAG can be combined to retrieve examples that are specific to the target input [19]. However, the standard approach for ICL and RAG uses text examples and inputs, whereas RRG is an image to text task.
@@
While general domain vision-language FMs are improving even on medical vision-language tasks [68, 48], there are an increasing number of medical modality specific FMs which may offer better performance for tasks over these modalities [51, 78, 80, 23, 74, 53, 10, 46, 4, 2, 21, 5, 81]. It is an open area of research on how to compose together these FMs which were not necessarily jointly trained [9, 45]. Furthermore, the composition or adaptation method depends on the task. For image captioning, text generation must generally be informed by the image features. We argue that the image features need not be high-dimensional latent or token embeddings, as in the current paradigm of multimodal LLMs.
Comparison LaB- RAG RAG SFT
@@
We propose LaB-RAG, Label Boosted Retrieval Augmented Generation, a method for image captioning via RAG and ICL without the need to train any deep learning models. We study LaB-RAG in the context of RRG. Fig. 1 shows a visual overview of LaB-RAG compared to standard RAG. Tab. 1 includes a high-level conceptual comparison of LaB-RAG to standard RAG and SFT methods for RRG.
Main Contributions:
  • A generalizable framework to use classical machine learning (ML) to boost LLM performance. By training simple ML models, e.g. logistic regression, to derive categorical labels from images, we propose to use the labels to inform RAG retrieval and to use the label names as textual image descriptors input to LLMs. This adapts from frozen pretrained vision models to frozen pretrained LLMs, only ever training classical models.
  • State of the art (SOTA) performance for RRG. We show that LaB-RAG for RRG can achieve SOTA when compared with RRG models from the literature. We further experiment with and analyze effects of different components of LaB-RAG.
  • Critique of a commonly used RRG metric. We critique the use of a label-based RRG metric commonly used in the literature, demonstrating that we can inflate its results by using its labels in LaB-RAG.
2 Related Work
Figure 2: LaB-RAG inference for RRG. Symbols correspond to those in Alg. 1.
@@ 2 Related Work
[…]
2:retrieval studies with image-label-text tuples
$S\leftarrow\{(i_{x},l_{x},t_{x})\}$
@@
[…]
6:compute retrieval embeddings
$E\leftarrow[...,\vec{e}_{x},...]^{\intercal}$,  $\vec{e}_{x}\leftarrow\theta_{img}(i_{x})$
7:binarize retrieval labels $B\leftarrow\{b_{x}\leftarrow l_{x}=1\}$
8:train image classification model $\psi_{img}(E)\rightarrow B$
[…]
10:compute image cosine similarity
[…]
11:sort studies by similarity $\vec{r}\leftarrow[...,s_{x},...]$,  $d_{x}\geq d_{x+1}$
12:Label Filter: filter or rerank $s_{x}\in\vec{r}$ by comparing
[…]
@@
[…]
As interest in RRG has steadily increased [69], there are an abundance of available RRG models from the literature trained to generate reports over CXRs. Additionally, the recent BioNLP workshop at ACL 2024 hosted a shared task on RRG [15] where several new models were presented. These published methods can be categorized by the report sections they generate, the “Findings” [73], “Impression” [18, 64, 28, 54, 65], both independently [12, 56], or both jointly [72]. Models from the literature can be further divided by the method for generation, either by a generative LLM conditioned only on image features [56, 12, 73, 54] or by text retrieval and processing [18, 64, 28, 54, 72, 65].
There are two primary ways by which LLMs are fine-tuned to generate the report based on input CXR embeddings. CXRMate [56] is trained with image embeddings input via cross attention [76, 8, 44]. CheXagent [12] and RGRG [73] are instead trained with image embeddings prepended as input tokens before the report, adapting the image embeddings into token embedding space.
The following retrieval-based methods use cross-modal image-to-text retrieval, requiring training of a retrieval model with a joint embedding space for CXRs and their corresponding reports. CXR-RePaiR-Gen [65] leverages CXR-ReDonE [64] for its cross-modal retrieval models and otherwise is the closest implementation of a standard RAG pipeline. FactMM-RAG [72] also employs RAG for inference, but trains its own retrieval model using RadGraph [27] labels to inform the joint embedding space. This is broadly similar to the training of X-REM’s [28] retrieval model, however X-REM uses CheXbert [70] labels. X-REM outputs a concatenation of retrieved text as the final report. CXR-RePaiR [18] also uses concatenation of retrieved text for its final output, however its retrieval model is trained via the basic CLIP [62] method. CXR-ReDonE [64] is the same as CXR-RePaiR except the training/retrieval data was cleaned to remove “priors” indicating a previous CXR [63].
The most closely related method compared to LaB-RAG is Pragmatic Retrieval/Llama [54]. Like LaB-RAG, Pragmatic derives categorical labels directly from the CXR. However, Pragmatic trains an end-to-end ResNet50 [22] model, whereas LaB-RAG uses simpler logistic classifiers trained over extracted image embeddings. Additionally, Pragmatic requires the report’s “Indication”, the clinical reason and motivation for the imaging study. While LaB-RAG uses both image embedding similarity and label matching for retrieval, Pragmatic Retrieval only uses label matching of image and indication for retrieval of report text. Pragmatic Retrieval then concatenates label-retrieved text as the final report. Pragmatic Llama does no retrieval, instead training an LLM to generate the report given the indication text and the positive image labels as text. LaB-RAG also uses labels as textual image descriptors but relies on ICL and model instructing, and thus does not require model fine-tuning.
3 LaB-RAG Framework
Figure 3: LaB-RAG does better than other retrieval methods (CXR-RePaiR/ReDonE, X-REM) and no different than SFT methods on “Impression” generation (CheXagent, CXRMate) and comparable to “Findings” generation measured by F1CheXbert (RGRG, CheXagent, CXRMate). See Sec. 5.3 for analysis of F1RadGraph results for “Findings” generation.
LaB-RAG uses a label boosted RAG algorithm with ICL to do image captioning. LaB-RAG retrieves paired example text using image embedding similarity. Retrieved texts are then fed into a general domain LLM with strong instruction following and natural language comprehension capabilities. We improve upon standard RAG by incorporating predicted categorical labels into both the example retrieval and text augmentation steps. The overview of LaB-RAG applied to RRG is presented in Fig. 2 with its high-level pseudocode described in Alg. 1. We evaluate our generation results against ground truth radiology reports using natural language metrics and clinical radiology language metrics. We compare our method against models from literature and against variations of our method.
3.1 Image, Text, and Label Dataset
For our study, we use chest X-rays, radiology reports, CheXpert [25] labels, metadata, and data splits from MIMIC-CXR v2.1.0 and its related datasets. [36, 32, 35, 33]. The dataset is split at the patient level. The final number of samples used in each of our experiments depends on the availability of all required data modalities. Further details on the dataset are provided in Sec. S2.
Because the radiology reports are written at the study level and the CheXpert [25] labeler derives labels from the report, the categorical labels are also defined per study. In order of preference, CheXpert labels are extracted from the “Impression”, “Findings”, or last paragraph sections of the report [32]. CheXpert provides 14 labels where each label describes a radiological observation, including “No Finding”. Each label gets a value of 1 (positive), 0 (negative), -1 (uncertain), or null (unmentioned). A report receives a multilabel classification where any combination of labels and values can be assigned. It is thus possible to have a study with no positive labels; in such cases, we assign a positive “Other” label which is negative in all other instances. Our final labels are sets of 15 labels per study.
3.2 Image Feature Extraction
We compute image feature embeddings to train our image classifier and to enable retrieval of visually similar images and their associated text. For CXR data, we use the pretrained X-ray image encoder of BioViL-T [2] to extract image features. We select BioViL-T as our image encoder because it was trained over MIMIC-CXR, it achieved impressive performance on tasks via linear probing, and because the hi-ml-multimodal v0.2.2 python package made it simple to integrate the model into our code. While the joint vision-language training may have helped BioViL-T learn more meaningful embeddings for downstream tasks, LaB-RAG is agnostic to the method by which the image encoder was trained. Specifically, we extract X-ray embeddings with the frozen image-encoder of BioViL-T. For image-based retrieval, we use projected 128d embeddings; for training our image classifier, we use unprojected 512d embeddings.

### v2

@@ Abstract
In the current paradigm of image captioning, deep learning models are trained to generate text from image embeddings of latent features. We challenge the assumption that fine-tuning of large, bespoke models is required to improve model generation accuracy. Here we propose Label Boosted Retrieval Augmented Generation (LaB-RAG), a small-model-based approach to image captioning that leverages image descriptors in the form of categorical labels to boost standard retrieval augmented generation (RAG) with pretrained large language models (LLMs). We study our method in the context of radiology report generation (RRG) over MIMIC-CXR and CheXpert Plus. We argue that simple classification models combined with zero-shot embeddings can effectively transform X-rays into text-space as radiology-specific labels. In combination with standard RAG, we show that these derived text labels can be used with general-domain LLMs to generate radiology reports. Without ever training our generative language model or image embedding models specifically for the task, and without ever directly “showing” the LLM an X-ray, we demonstrate that LaB-RAG achieves better results across natural language and radiology language metrics compared with other retrieval-based RRG methods, while attaining competitive results compared to other fine-tuned vision-language RRG models. We further conduct extensive ablation experiments to better understand the components of LaB-RAG. Our results suggest broader compatibility and synergy with fine-tuned methods to further enhance RRG performance. Our code is available at: https://github.com/uc-cdis/label-boosted-RAG-for-RRG.
@@ 1 Introduction
Radiology reports are free-text natural language notes describing the observations seen in radiological images, such as X-rays, CT scans, or MRI scans. These reports are written by board-certified radiologists, highly specialized doctors (Rosenkrantz et al., 2020) who are in worsening short supply (Kumar et al., 2020; Christensen et al., 2023; Rimmer, 2017; Ismail et al., 2024). Motivated by the popularization of large language models (LLMs), there has been an increasing interest in AI tools to help bridge the radiologist shortage gap (Hosny et al., 2018; Najjar, 2023; Kelly et al., 2022).
@@
Radiology report generation (RRG) is the task of automatically generating these reports given the images (Sloan et al., 2024; Messina et al., 2022). While RRG can be applied to any radiological imaging modality, the most frequent modality studied in the literature is chest X-rays (CXRs). This is evidenced by the large number of publicly available paired CXR-report datasets (Sloan et al., 2024), such as MIMIC-CXR (Johnson et al., 2019), CheXpert Plus (Chambon et al., 2024), and others (Demner-Fushman et al., 2016; Bustos et al., 2020; Vayá et al., 2020; Nguyen et al., 2022). For CXRs, RRG is typically formulated as generating the “Findings” or “Impression” section of the report (Messina et al., 2022). Conceptually, the findings section describes all positive or negative observations seen in the X-ray, while the impression section summarizes and interprets those findings with recommendations for clinical diagnosis (ESR, 2011).
Figure 1: Overview of LaB-RAG for RRG compared to standard RAG.
At its core, RRG is a form of image captioning, specialized for medical imaging (Stefanini et al., 2022). In recent years, LLMs have demonstrated impressive performance across medicine (Nori et al., 2023; Chen et al., 2023; Saab et al., 2024). For other medical vision-language tasks, efforts typically involve training models for the target task (Beddiar et al., 2023; Ayesha et al., 2021), particularly when starting with an out-of-domain foundation model (FM) (Bommasani et al., 2021).
@@
Model adaptation classically involves model training via supervised fine tuning (SFT). Yet, SFT of FMs is becoming increasingly difficult as models become larger, requiring compute resources greater than consumer-grade workstations can provide (Tuggener et al., 2024). While parameter-efficient fine tuning (PEFT) methods have demonstrated competitive results (Ding et al., 2023), a form of LLM adaptation that does not require model training is in-context learning (ICL) (Dong et al., 2022). The goal of ICL is to have an LLM infer the target output using examples of input-output pairs given jointly at inference time with the target input (Min et al., 2022). Related to ICL is retrieval augmented generation (RAG), a framework for providing additional context to an LLM prompt by retrieving documents related to the input query (Lewis et al., 2020). ICL and RAG can be combined to retrieve examples that are specific to the target input (Gao et al., 2023). However, ICL and RAG applied to the image to text task of RRG requires considering model modality.
@@
While general domain vision-language FMs are improving even on medical vision-language tasks (Saab et al., 2024; Meta, 2024), there are an increasing number of medical modality specific FMs which have demonstrated stronger results (Moor et al., 2023; Wornow et al., 2023; Zhang & Metaxas, 2024; He et al., 2024; Thieme et al., 2023; Neidlinger et al., 2024; Chen et al., 2024a; Lu et al., 2024; Boecking et al., 2022; Bannur et al., 2023; Gu et al., 2021; Bolton et al., 2024; Zhang et al., 2023). It is an open area of research on how to compose together multiple FMs which were not necessarily jointly trained (Chen et al., 2022; Lin et al., 2024). Such composition depends on the task; in image captioning, the image must inform the text generation. We argue that the image features need not solely be high-dimensional latent embeddings, as with modern multimodal LLMs.
We propose LaB-RAG, Label Boosted Retrieval Augmented Generation, a method for image captioning which improves upon RAG and ICL. We study LaB-RAG in the context of RRG. Figure 1 and Table 1 present conceptual comparisons of LaB-RAG versus standard RAG and SFT methods.
Our main contributions are as follows:
  • A modular framework for image captioning using rich embeddings coupled with small to mid-scale models. By training simple machine learning (ML) models (e.g. logistic regression) over zero-shot image embeddings to derive categorical labels, we use the labels to filter RAG retrieved text and to contextualize ICL examples. LaB-RAG composes frozen, disjoint vision and language models at the low cost of training classical ML models agnostic to the downstream task.
  • State of the art performance for RRG. On clinical radiology metrics, we demonstrate that LaB-RAG for RRG beats other retrieval based models, and we show that LaB-RAG achieves state of the art performance when compared with SFT models from the literature.
  • A novel framework complementary to training methods that improve models. Through extensive ablation experiments of LaB-RAG over the two largest public CXR datasets, we better understand the potential synergy of alternate modeling choices. Our results suggest that LaB-RAG is complementary to SFT and other methods.
2 Related Work
Comparison
LaB-RAG
RAG SFT
@@ 2 Related Work
As interest in RRG has steadily increased (Sloan et al., 2024), there are an abundance of available RRG models from the literature trained to generate reports over CXRs. Additionally, the recent BioNLP workshop at ACL 2024 hosted a shared task on RRG (Demner-Fushman et al., 2024) where several new models were presented. These published methods can be categorized by the report sections they generate, the “Findings” (Tanida et al., 2023), “Impression” (Endo et al., 2021; Ramesh et al., 2022b; Jeong et al., 2024; Nguyen et al., 2023; Ranjit et al., 2023), both independently (Chen et al., 2024b; Nicolson et al., 2024a), or both jointly (Sun et al., 2024). Models from the literature can be further divided by the method for generation, either by a trained LLM conditioned on high-dimensional image embeddings (Nicolson et al., 2024a; Chen et al., 2024b; Tanida et al., 2023; Nguyen et al., 2023) or by text retrieval and processing (Endo et al., 2021; Ramesh et al., 2022b; Jeong et al., 2024; Nguyen et al., 2023; Sun et al., 2024; Ranjit et al., 2023).
There are two primary ways by which LLMs are fine-tuned to generate the report based on input CXR embeddings. CXRMate (Nicolson et al., 2024a) is trained with image embeddings input via cross attention (Vaswani et al., 2017; Chen et al., 2021; Lin et al., 2022). CheXagent (Chen et al., 2024b) and RGRG (Tanida et al., 2023) are instead trained with image embeddings prepended as input tokens before the report, adapting the image embeddings into token embedding space.
The following retrieval-based methods use cross-modal image-to-text retrieval, requiring training of a retrieval model with a joint embedding space for CXRs and their corresponding reports (Zhang et al., 2022). CXR-RePaiR-Gen (Ranjit et al., 2023) leverages CXR-ReDonE (Ramesh et al., 2022b) for its cross-modal retrieval models and otherwise is the closest implementation of a standard RAG pipeline. FactMM-RAG (Sun et al., 2024) also employs RAG for inference, but trains its own retrieval model using RadGraph (Jain et al., 2021) labels to inform the joint embedding space. This is broadly similar to the training of X-REM’s (Jeong et al., 2024) retrieval model, however X-REM uses CheXbert (Smit et al., 2020) labels. X-REM outputs a concatenation of retrieved text as the final report. CXR-RePaiR (Endo et al., 2021) also uses concatenation of retrieved text for its final output, however its retrieval model is trained via the basic CLIP (Radford et al., 2021) method. CXR-ReDonE (Ramesh et al., 2022b) is the same as CXR-RePaiR except the training/retrieval data was cleaned to remove “priors” indicating a previous CXR (Ramesh et al., 2022a).
The most closely related method compared to LaB-RAG is Pragmatic Retrieval/Llama (Nguyen et al., 2023). Like LaB-RAG, Pragmatic derives categorical labels directly from the CXR. However, Pragmatic trains an end-to-end ResNet50 (He et al., 2016) model, whereas LaB-RAG uses simpler logistic classifiers trained over extracted image embeddings. Additionally, Pragmatic requires the report’s “Indication”, the clinical motivation for the imaging study. While LaB-RAG uses both image embedding similarity and label matching for retrieval, Pragmatic Retrieval only uses label matching of image and indication for retrieval of report text. Pragmatic Retrieval concatenates label-retrieved text as the final report. Pragmatic Llama does no retrieval, instead training an LLM to generate the report given the indication text and the positive image labels as text. LaB-RAG also uses labels as textual image descriptors but relies on ICL, and thus does not require LLM training.
3 LaB-RAG Framework
Figure 2: LaB-RAG inference for RRG. Symbols correspond to those in Algorithm 1.
LaB-RAG is a label boosted RAG algorithm with ICL to do image captioning. LaB-RAG retrieves paired example text using image embedding similarity. Retrieved texts are then fed into a general domain LLM with strong instruction following and natural language comprehension capabilities. We improve upon standard RAG by incorporating predicted categorical labels into both the example retrieval and text augmentation steps. The overview of LaB-RAG applied to RRG is presented in Figure 2 with its high-level pseudocode described in Algorithm 1. We study LaB-RAG on two CXR datasets, MIMIC-CXR (Johnson et al., 2019) and CheXpert Plus (Chambon et al., 2024). We conduct extensive experimental comparisons and evaluate our generated reports against ground truth reports using natural language and radiology-specific metrics.
3.1 Image, Text, and Label Data
For our study, we use chest X-rays, radiology report sections, metadata, and data splits from either MIMIC-CXR v2.1.0 (Johnson et al., 2024) or CheXpert Plus (Chambon et al., 2024). The datasets are split at the patient level. We extract categorical labels from the ground-truth radiology reports using either the CheXbert (Smit et al., 2020) or CheXpert (Irvin et al., 2019) labeler; we specifically extract labels from the report section we aim to generate, i.e. Findings or Impression. The final number of samples used in each of our experiments depends on the availability of all required data modalities. Further details on the datasets are provided in Section B.1.
As reports are written at the study level and the extracted labels are derived from the report, the categorical labels are also defined per study. Both labelers extract the same 14 label types, where each label describes an observation, including “No Finding”. Each label gets a value of 1 (positive), 0 (negative), -1 (uncertain), or null (unmentioned). Given the multilabel multiclass assignment, it is possible to have a study with no positive labels; in such cases, we assign a positive “Other” label which is negative in all other instances. Our final labels are sets of 15 labels per study.
3.2 Image Embeddings
We extract zero-shot, frozen image feature embeddings to enable retrieval of similar images with their associated text and to train our image classifier. For our experiments, we utilize two domain-adapted image models, BioViL-T (Bannur et al., 2023) and GLoRIA (Huang et al., 2021). To enable more fair comparisons, we select these models for their similarity in contrastive pretraining (Oord et al., 2018) (though LaB-RAG is training objective agnostic) and their reported high performance on linear probing tasks and embedding-based retrieval. Importantly, we use only use BioViL-T on MIMIC-CXR experiments and GLoRIA on CheXpert Plus, as we observed strong training dataset specificity in early preliminary experiments. We further compare both to ResNet50 (He et al., 2016) trained on ImageNet (Deng et al., 2009). See Section B.3 for further details.
@@ 3.2 Image Embeddings
1:inference image $i_{\mathit{inf}}$
2:retrieval studies with image-label-text tuples $S\leftarrow\{s_{x}\leftarrow(i_{x},l_{x},t_{x})\}$
@@
5:compute inference embedding $\vec{e}_{\mathit{inf}}\leftarrow\theta_{img}(i_{\mathit{inf}})$
6:compute retrieval embeddings $E\leftarrow\langle\vec{e}_{x}\mid\vec{e}_{x}=\theta_{img}(i_{x})\rangle^{\intercal}$
7:binarize retrieval labels $B\leftarrow\langle b_{x}\mid b_{x}=\mathbf{1}_{\{l_{x}=1\}}\rangle$
8:train image classification model $\psi_{img}\leftarrow\arg\min_{\psi}\mathcal{L}_{\text{BCE}}(\psi,E,B),\psi:E\to B$
9:infer inference image label $\hat{b}_{\mathit{inf}}\leftarrow\psi_{img}(\vec{e}_{\mathit{inf}})$
10:compute image cosine similarity $\vec{d}\leftarrow\langle d_{x}\mid d_{x}=(\vec{e}_{\mathit{inf}}\cdot\vec{e}_{x})~/~\lVert\vec{e}_{\mathit{inf}}\rVert\lVert\vec{e}_{x}\rVert\rangle$
11:sort studies by similarity $\vec{r}\leftarrow\langle s_{x}\mid d_{x}\geq d_{x+1}\rangle$
12:Label Filter: filter or rerank $s_{x}\in\vec{r}$ by comparing labels $\hat{b}_{\mathit{inf}}$ to $b_{x}\in B$ (e.g. filter $b_{x}=\hat{b}_{\mathit{inf}}$)
@@
14:prepare prompt $p_{\mathit{inf}}$ using retrieved text $t_{x}\in S_{k}$
15:Label Format: format $p_{\mathit{inf}}$ with labels $\hat{b}_{\mathit{inf}}$ and $l_{x}\in S_{k}$ (e.g. list positive labels $l=1$)
16:generate report $\hat{t}_{\mathit{inf}}\leftarrow\phi_{txt}(p_{\mathit{inf}})$
17:generated report $\hat{t}_{\mathit{inf}}$